A parallel plate capacitor with plate separation d is connected to a battery. The capacitor is fully charged to Q Coulombs and a voltage of V. (C is the capacitance and U is the stored energy.) Answer the following questions regarding the capacitor charged by à battery. For each statement below, select True or False. 8 5 8 2 5 After being disconnected from the battery, inserting a dielectric with k will decrease V. After being disconnected from the battery, increasing d decreases V After being disconnected from the battery, decreasing d increases U After being disconnected from the battery, decreasing d decreases C. With the capacitor connected to the battery, inserting a dielectric with K will decrease U Submit Answer Incorrect. Tries 1/99 Previous Tries
Answer #1

C=kEoA/d ,V=Q/C,U=1/2 CV^2

After being disconnected from the battery, inserting a dielectric with k will decrease V.---True

V decreases since V=Q/C and c=keo A/d so V is inversely proportional to k

After being disconnected from the battery, increasing d decreases V---False --it increases

After being disconnected from the battery, decreasing d increases u- true

After being disconnected from the battery, decreasing d decreases C--False it increases

With the capacitor connected to the battery, inserting a dielectric with k will decrease C--- False, it will increase